User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Meeting notes from 5th May 2016

Notes: Only hb 12m in operation over last two weeks

Chair/Presenting: Jesse/Ross

Attendees: Jesse, Ross, Liz, Jonny, Ellen, Lucas, Arwin, Lucia, Jim

Bryn, Tiege are unable to attend.


  • hb is behaving well except for the usual tsys errors associated with RFI and weather.
  • When noting high tsys in ERC please specify channels and approximate values. e.g., rather than “tsys high in both bands” you could say “High tsys (~170–250K) for chan 07,09,12,14
  • R1737 – Slow start to recording of each scan? (~50gb data loss)
  • RD1604 may19: chang'E observations
  • possible shift friday morning (Thanks Arwin)


  • Tsys issues with R1737 at hb seem to be a one off occurrence.
  • Both hb and yg need to be moved to continuous tsys measurements (When Jamie has a spare day or so)
/home/www/auscope/opswiki/data/pages/operations/agenda2016_05_06.txt · Last modified: 2016/05/05 02:27 by Jesse Swan